A Morgan Stanley Consumer Research Survey sampled men and women and asked each whether they preferred to drink plain bottled water or a sports drink such as Gatorade or Propel Fitness water. Suppose 200 men and 200 women participated in the study, and 280 reported they preferred plain bottled water. Of the group preferring a sports drink, 80 were men and 40 were women.

                    

(a)  What is the probability a person in the study preferred plain bottled water?

(b)What is the probability a person in the study preferred a sports drink?


 


1
Expert's answer
2021-09-13T00:30:22-0400

(a)

The following table shows the sample data.



The joint probability of the given data is as follows:



The probability that a person preferred plain bottled water is,

"P (B) = \frac{120+160}{400}= 0.7"

(b)

The probability that a person preferred sports drink is,

"P (S) = \frac{80+40}{400}=0.3"